OpenStudy (anonymous):

Solve the equation. Be sure to show your work. All final answers must be in simplest form. b+1/9=9

OpenStudy (anonymous):

b=9-1/9 b=17/9 simplest form

OpenStudy (anonymous):

what? since when is \[80+\frac{1}{9}=9\]?

OpenStudy (anonymous):

is question \[\frac{b+1}{9}=9\]? if so answer is 80. if it is\[b+\frac{1}{9}=9\] then answer is \[\frac{80}{9}\]

OpenStudy (anonymous):

b+ 1/9 =9!

OpenStudy (anonymous):

b=9-(1/9) b=( 81-1)/9 b= 80/9